OpenStudy (anonymous):

Suppose you roll an eight-sided die two times hoping to get two numbers whose sum is greater than ten. What is the sample space? How many favorable outcomes are there?

OpenStudy (anonymous):

min number on each dice is 1 and max would be 8 as its 8 sided, so sum greater than 10 possible events are : {(2,8) ( 3,7) , (3,8) ,(4,6) ,(4,7) (4,8) (5,5) (5,6) (5,7) (5,8) (6,6) (6,7) (6,8) (7,7) (7,8) (8,8)} that is our favourable outcomes sample space. the total sample space would contain 8*8 = 64 pairs so the probability as we know is favourable / total events so we have p = 16/64 = 1/4 = 0.25

OpenStudy (anonymous):

pardon, my answer is correct if he asked sum greater or equal to 10, for ur problem , u have delete some favourable cases , can u tell what are they?